Trace the curve y= cube root of x^2 -1 and state all the properties you use to trace it.
1
Expert's answer
2020-03-23T10:03:50-0400

"y=\sqrt[3]{x^2-1}"

1) The domain "x\in(-\infty;\infty)" .

 2) Evenness, oddity

"y(-x)=\sqrt[3]{(-x)^2-1}=\sqrt[3]{x^2-1}=y(x)"

graph is symmetric about the axis "Oy".

3)  Continuity

the function is continuous "x\in (-\infty;\infty)"

 4) Periodicity

the function is not periodic

 5) Intersection points with coordinate axes

"Ox: y=0\\
\sqrt[3]{x^2-1}=0\\
x^2-1=0\\
x^2=1\\
x_1=1, x_2=-1\\
A(1,0), B(-1,0)\\
Oy:x=0\\
y=\sqrt[3]{0^2-1}=\sqrt[3]{-1}=-1\\
C(0,-1)"

6) Let's check the extremum

"y'=((x^2-1)^\frac{1}{3})'=\frac{2}{3}\cdot x\cdot(x^2-1)^{-\frac{2}{3}}\\
y'=0\\
\frac{2}{3}\cdot x\cdot(x^2-1)^{-\frac{2}{3}}=0\\
x_1=0, x_2=1, x_3=-1\\
x\in(-\infty;-1)\cup(-1;0], y\searrow\\
x\in [0;1)\cup(1,\infty),y\nearrow\\
x=0-min\\
y_{min}=-1"

7) Convex intervals

"y''=\frac{2}{3}(x^2-1)^{-\frac{2}{3}}-\frac{8}{9}\cdot x^2\cdot(x^2-1)^{-\frac{5}{3}}\\
y''=0\\
\frac{2}{3}(x^2-1)^{-\frac{2}{3}}-\frac{8}{9}\cdot x^2\cdot(x^2-1)^{-\frac{5}{3}}=0\\
(x^2-1)^{-\frac{5}{3}}(x^2-1-\frac{4}{3}x^2)=0\\
x_1=-1,x_2=1, \frac{1}{3}x^2=-1\\
x\in (-\infty;-1)\cup(1,\infty)"

since the second derivative is at this interval

"y''<0" then

"y" is convex upward "\cap"

"x\in (-1,1)"

since the second derivative is at this interval

"y''>0" then

"y" is convex down "\cup"

"x=1, x=-1" are inflection points

8) Asymptotes

 there are no vertical asymptotes

 find oblique, horizontal asymptote "y=kx+b"

"k=\lim\limits_{ x\to\infty}\frac{f(x)}{x}=\lim\limits_{ x\to\infty}\frac{\sqrt[3]{x^2-1}}{x}=0\\
b=\lim\limits_{ x\to\infty}(f(x)-kx)=\lim\limits_{ x\to\infty}(\sqrt[3]{x^2-1}-0)=\infty"

there are no oblique, horizontaltical asymptote "y=kx+b"

9)

"\lim\limits_{x\to -\infty}\sqrt[3]{x^2-1}=+\infty\\
\lim\limits_{x\to\infty}\sqrt[3]{x^2-1}=+\infty"